OpenStudy (anonymous):

Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (please explain to me. I'm not great with systems of equations.)

OpenStudy (anonymous):

You know you have 46 coins, and you have dimes and quarters. The total of the dimes and quarters needs to add up to 7 dollars. The systems of equations should be: 0.1d + 0.25q = 7 d + q = 46 does this help? (:

OpenStudy (anonymous):

should have mentioned this as well: the first equation shows that 0.1d (the total amount from the dimes) plus 0.25q (the total amount from te quarters) should equal 7 dollars. the second equation shows that the number of dimes and quarters must add up to 46 coins. ~ After this, you shold solve for either d or q. Let's say you were to solve for q, quarters. You multiply the second equation by -0.1 (negative, because we need to cancel out the d variable). -0.1 (d + q = 46) --> -0.1d - 0.1q = -4.6 So our equations are now: 0.1d + .25q = 7 -0.1d - 0.1q = -4.6 ~ Now add both of the equations. You can cancel 0.1d and -0.1d, leaving you with: .24q = 2.4 divide both sides by .24: q = 10 ~ So there are 10 quarters. With this information, you can solve for d, the number of dimes Hope this helped!! C;

OpenStudy (anonymous):

It's a bit hard to understand the work but I think I kind of get it.